suppose that no of accident occuring on high way is possion with random vairables lamda=3


1
Expert's answer
2022-06-07T09:02:13-0400

a.


"P(X\ge 3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dfrac{e^{-3}(3)^0}{0!}-\dfrac{e^{-3}(3)^1}{1!}"

"-\dfrac{e^{-3}(3)^2}{2!}=0.576810"

b.


"P(X\ge3|X\ge1)=\dfrac{P(X\ge3\cap X\ge 1)}{P(X\ge 1)}"

"=\dfrac{P(X\ge3)}{P(X\ge 1)}=\dfrac{1-\dfrac{e^{-3}(3)^0}{0!}-\dfrac{e^{-3}(3)^1}{1!}-\dfrac{e^{-3}(3)^2}{2!}}{1-\dfrac{e^{-3}(3)^0}{0!}-\dfrac{e^{-3}(3)^1}{1!}}"

"=\dfrac{0.576610}{0.800852}=0.720246"